OpenStudy (anonymous):

−4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution awnser ?

OpenStudy (anonymous):

x=-1 or x=-9

jigglypuff314 (jigglypuff314):

to solve for x for −4|x + 5| = −16 first want to isolate the absolute values so /-4 /-4 divide both sides by -4 |x + 5| = 4 now separate into two equations x + 5 = 4 and -(x + 5) = 4 and solve for x so -5 -5 - x - 5 = 4 x = 4 - 5 x = - 5 - 4 then simplify

OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution i need help with this onetoo

jigglypuff314 (jigglypuff314):

to solve for x for |x − 2| + 10 = 12 first isolate the absolute values -10 -10 |x - 2| = 2 then separate into two equations x - 2 = 2 and -(x - 2) = 2 then solve for x so +2 +2 - x + 2 = 2 x = 2 + 2 x = 2 - 2 then simplify

OpenStudy (anonymous):

x=0 or x=4